# Large Language Models for Education: A Survey and Outlook

### v1

@@ 1. Introduction
As large language models (LLMs) have become a powerful paradigm in different areas (Fan et al., 2023b; Zeng et al., 2023; Jin et al., 2024; Chen et al., 2023a), they also achieved state-of-the-art performances in multiple educational scenarios (Li et al., 2023a; Kasneci et al., 2023; Yan et al., 2024). Existing work has found that LLMs can achieve student-level performance on standardized tests (OpenAI, 2023) in a variety of mathematics subjects (e.g., physics, computer science) on both multiple-choice and free-response problems. In addition, empirical studies have shown that LLMs can serve as a writing or reading assistant for education (Malinka et al., 2023; Susnjak, 2022). A recent study (Susnjak, 2022) reveals that ChatGPT is capable of generating logically consistent answers across disciplines, balancing both depth and breadth. Another quantitative analysis (Malinka et al., 2023) shows that students using ChatGPT (by keeping or refining the results from LLMs as their own answers) perform better than average students in some courses from the computer security field. Recently, several perspective papers (Tan et al., 2023; Kamalov and Gurrib, 2023) also explore various application scenarios of LLMs in classroom teaching, such as teacher-student collaboration, personalized learning, and assessment automation. However, the application of LLMs in education may lead to a series of practical issues, e.g., plagiarism, potential bias in AI-generated content, overreliance on LLMs, and inequitable access for non-English speaking individuals (Kasneci et al., 2023).
Despite the rapid developments and promising results of LLM for education, the previous literature has not systematically summarized LLMs for education from a technological perspective. To bridge this gap, this survey aims to provide a comprehensive technological review of LLMs for education, which provides a novel technology-centric taxonomy and summary of existing publicly available datasets and benchmarks. Furthermore, we summarize current challenges as well as further research opportunities, in order to foster innovations and understanding in the dynamic and ever-evolving landscape of LLMs for education. In summary, our contributions lie in the following three main parts:
@@
3, Current challenges and future research directions We discuss current risks and challenges, as well as highlight future research opportunities and directions, urging researchers to delve deeper into this exciting area.
@@ 2.2. Study Assisting
Providing students with timely learning support has been well recognized as a crucial factor in improving students’ engagement and learning efficiency during their independent studies (Dewhurst et al., 2000). Due to the limitation of prior algorithms in generating fixed-form responses, many of the existing study-assisting approaches face poor generalization challenges while being implemented in real-world scenarios (König et al., 2023). Fortunately, the appearance of LLMs brings revolutionary changes to this field. By taking advantage of instruct fine-tuned LLMs (Ouyang et al., 2022) in generating human-like responses, many recent LLM-based study assisting works have presented appealing results in providing real-time supports to students in solving blocking questions, correcting students’ errors, and providing explanations or hints to their confusions.
@@ 2.2.1. Question Solving (QS)
Contributing to the large-scale parameter size of LLMs and the enormous sized and diverse web corpus used during the pre-training phase, LLMs have been proven to be a powerful question zero-shot solver to questions spread from a wide spread of subjects, including math (Yuan et al., 2023; Wu et al., 2023c), law (Bommarito II and Katz, 2022; Cui et al., 2023), medicine (Thirunavukarasu et al., 2023; Liévin et al., 2023), finance (Wu et al., 2023b; Yang et al., 2023), programming (Kazemitabaar et al., 2023; Savelka et al., 2023), language understands(Zhang et al., 2024; Achiam et al., 2023). In addition, to further improve LLM’s problem-solving performance while facing complicated questions, a variety of studies have been actively proposed. For example, Wei et al. (2022) propose the Chain-of-Thought (CoT) prompting method, which guides LLMs to solve a challenging problem by decomposing it into simpler sequential steps. Other works (Sun, 2023; Wang et al., 2023) exploit the strong in-context learning ability of LLMs and propose advanced few-shot demonstration-selection algorithms to improve LLM’s problem-solving performance to general questions. Chen et al. (2022) and Gao et al. (2023) leverage external programming tools to avoid calculation errors introduced during the textual problem solving process of raw LLMs. Wu et al. (2023a) regard chat-optimized LLMs as powerful agents and design a multi-agent conversation to solve those complicated questions through a collaborative process. Cobbe et al. (2021) and Zhou et al. (2024b) propose the external verifier module to rectify intermediate errors during generation, which improves LLM problem solving performance on challenging math questions. Overall, with the proposition of all these novel designs, the use of LLMs for question solving has achieved impressive progress. Furthermore, students can find high-quality answers to their blocking questions in a timely manner.
@@ 3. Dataset and Benchmark
Datasets and benchmarks for educational applications vary widely in scope and purpose, targeting different aspects of the educational process, such as student performance data (Ray et al., 2003), text and resource databases (Brooke et al., 2015), online learning data (Ruipérez-Valiente et al., 2022), language learning database (Tiedemann, 2020), education game data (Liu et al., 2020), demographic and socioeconomic data (Cooper et al., 2020), learning management system (LMS) data (Conijn et al., 2016), special education and needs data (Morningstar et al., 2017). LLMs revolutionized the field of natural language processing (NLP) by enabling a wide range of text-rich downstream tasks, which leverage the extensive knowledge and linguistic understanding embedded within LLMs to perform specific functions requiring comprehension, generation, or text transformation. Therefore, many datasets and benchmarks are constructed for text-rich educational downstream tasks.
Table 1 summarizes some commonly used publicly available datasets and benchmarks for evaluating LLMs on education applications111While there exist various educational applications, including those that assist with confusion helper, material creation, knowledge tracking, and content personalizing, we did not discuss them in this section due to the absence of publicly accessible datasets.. The Majority of datasets and benchmarks lie in the tasks of question-solving (QS), error correction (EC), question generation (QG), and automatic grading (AG), which covers the use cases that benefit different users, subjects, levels, and languages. Among all these datasets, some mainly benefit the student, while the others help the teacher.
@@
The question-solving ones (Cobbe et al., 2021; Hendrycks et al., 2020; Huang et al., 2016; Wang et al., 2017; Zhao et al., 2020; Amini et al., 2019; Miao et al., 2021; Lu et al., 2021b; Kim et al., 2018; Lu et al., 2021a; Chen et al., 2023b), take a significant amount as a prevalent task for both Education and NLP fields. In particular, many datasets (Cobbe et al., 2021; Hendrycks et al., 2020; Huang et al., 2016; Wang et al., 2017; Zhao et al., 2020; Amini et al., 2019; Miao et al., 2021; Lu et al., 2021b) are constructed for math question-solving, aiming to provide an abstract expression from a narrative description. Some datasets also take the images (Miao et al., 2021; Kim et al., 2018; Lu et al., 2021a; Kembhavi et al., 2016) and tables (Lu et al., 2021b) into consideration. On the other hand, another bunch of datasets and benchmarks (Kim et al., 2018; Kembhavi et al., 2016; Chen et al., 2023b), are constructed for science textbook question-solving, which requires a comprehensive understanding of the textbook and provide the answer corresponding to the key information in the question. There are also large amounts of datasets and benchmarks constructed for error correction. They are used for foreign language training (Rothe et al., 2021; Ng et al., 2014; Bryant et al., 2019; Tseng et al., 2015; Zhao et al., 2022; Xu et al., 2022b; Du et al., 2023; Náplava et al., 2022; Rozovskaya and Roth, 2019; Grundkiewicz and Junczys-Dowmunt, 2014; Davidson et al., 2020; Syvokon and Nahorna, 2021; Cotet et al., 2020) and computer science programming language training (Just et al., 2014; Le Goues et al., 2015; Lin et al., 2017; Tufano et al., 2019; Li et al., 2022; Guo et al., 2024). The foreign language training datasets and benchmarks contains grammatical errors and spelling errors that are needed to be identified and corrected. The programming training datasets and benchmarks include several code bugs that require sufficient coding understanding for proper correction.
On the other hand, there are several datasets and benchmarks for teacher-assisting tasks. (Welbl et al., 2017; Lai et al., 2017; Xu et al., 2022a; Chen et al., 2018; Gong et al., 2022; Hadifar et al., 2023; Liang et al., 2018; Bitew et al., 2022) are constructed for question construction task that aims to evaluate the ability of the LLM for generating the educational questions from a given context. (Yang et al., 2023; Tigina et al., 2023; Blanchard et al., 2013; Stab and Gurevych, 2014) are built for automatic grading the student assignments.
@@
4. Risks and Potential Challenges
4.1. Fairness and Inclusiveness
4.2. Reliability and Safety
4.3. Transparency and Accountability
LLM, by design, runs as a black-box mechanism, so it comes with transparency and accountability concerns. Milano et al. (Milano et al., 2023) and BaHammam et al. (BaHammam et al., 2023) raise several challenges about LLMs to higher education, including plagiarism, inaccurate reporting, cheating in exams, as well as several other operational, financial, pedagogical issues. As a further thought of students using generative AI in homework or exams, Macneil et al. (MacNeil et al., 2024) discuss the respective impact to the traditional assessment methods, and think we educators should come up with new assessment framework to take into account the usage of Chat-GPT-like tools. Zhou et al. (Zhou et al., 2024a) brought up academic integrity ethical concerns that specifically confused teachers and students and calls for rethinking of policy-making. As specific measures, Gao et al. (Gao et al., 2024a) introduce a novel concept called mixcase, representing a hybrid text form involving both machine-generated and human-generated and developed detectors that can distinguish human and machine texts. To tackle LLM ethical concerns in intellectual property violation, Huang et al. (Huang and Chang, 2023) propose incorporating citation while training LLMs, which could help enhance content transparency and verifiability. Finlayson, et al. (Finlayson et al., 2024) developed a systematic framework to efficiently discover the LLM’s hidden size, obtaining full-vocabulary outputs, detecting and disambiguating different model updates, which could help users hold providers accountable by tracking model changes, thus enhancing the accountability.
4.4. Privacy and Security
4.5. Overly Dependency on LLMs
Given the impressive performance of LLM generative ability, there is great concern that students blindly rely on LLMs for most of their work, leaving their ability to think independently disappearing. Milano et al. (Milano et al., 2023) discuss the problem of overreliance caused by Chat-GPT-like applications that students may use to compose essays and academic publications without improving their writing skills, which is essential to cultivate critical thinking. The concern might have even more impact on foreign-language students or students who are educationally disadvantaged while placing less emphasis on learning how to craft well-written texts. Krupp et al. (Krupp et al., 2023) discuss the challenges of overreliance on the implications of LLMs in education and propose some moderated approaches to mitigate such effects. Similarly, Zuber et al. (Zuber and Gogoll, 2023) discuss the risk that overreliance can bring to democracy and suggest cultivating thinking skills in children, fostering coherent thought formulation, and distinguishing between machine-generated output and genuine. They argue that LLMs should be used to augment but not substitute human thinking capacities. Adewumi et al. (Adewumi et al., 2023) also present scenarios that students tend to rely on LLMs for essay writing rather than writing their own, and demonstrates that using a probing-chain-of-thought tool could substantially stimulate critical thinking in the accompany with LLMs.
5. Future Directions
Here, we discuss the future opportunities for LLMs in education.
5.1. Multimodal Learning Analytics
The evolution of LLMs extends beyond textual analysis, venturing into the realm of multimodal learning analytics. This approach takes advantage of various types of data, including text, audio, video, and kinesthetic interactions, to provide a holistic understanding of the learning process. Future research could focus on developing LLMs capable of interpreting and integrating these varied data sources, offering more nuanced insights into student engagement, comprehension, and learning styles. Such advances could pave the way for highly personalized and adaptive learning experiences that are tailored to the unique needs of each student.
The following are some potential impacts of multi-modal learning analytics in education. First, by understanding student interactions across multiple modalities, LLMs can offer highly personalized feedback and learning pathways. Second, LLMs can facilitate a more holistic approach to evaluation than traditional assessments, incorporating not just what the student has learned, but how it interacts with the material. Third, multi-modal data can reveal patterns indicative of learning difficulties much earlier than conventional methods, which can help with the early detection of learning challenges.
5.2. Multilingual LLMs for Education
The expansion of LLMs into multilingual and non-English domains represents a crucial evolution in the educational landscape. Traditionally, English has dominated the digital and educational technology spaces, limiting access and equity for non-English speakers. Research could focus on developing robust models that not only translate, but also understand cultural nuances, colloquial expressions, and regional educational standards. This would help learners around the world to benefit from LLMs in their native languages and significantly improve equity and inclusion in global education.
More specifically, there are some possible impacts on the development of multilingual LLMs for education. First, by providing educational content and tools in a variety of languages, multilingual LLMs can significantly increase access to quality education. This is especially important for regions where educational resources are scarce or where there is a strong preference for instruction in the local language. Second, multilingual LLMs could offer opportunities to integrate cultural context and nuances into educational content, making learning more relevant and engaging for students. This cultural sensitivity can improve comprehension and retention, fostering a deeper connection with the material. Third, multilingual LLMs could serve as powerful tools for language learning, offering nuanced context-aware language models that can adapt to different proficiency levels.
5.3. Edge Computing and Efficiency
5.4. Efficient Training of Specialized Models
5.5. Ethical and Privacy Considerations
6. Conclusion

### v2

@@ 1. Introduction
As large language models (LLMs) have become a powerful paradigm in different areas (Fan et al., 2023b; Zeng et al., 2023; Jin et al., 2024; Chen et al., 2023a), they also achieved state-of-the-art performances in multiple educational scenarios (Li et al., 2023a; Kasneci et al., 2023; Yan et al., 2024). Existing work has found that LLMs can achieve student-level performance on standardized tests (OpenAI, 2023) in a variety of mathematics subjects (e.g., physics, computer science) on both multiple-choice and free-response problems. In addition, empirical studies have shown that LLMs can serve as a writing or reading assistant for education (Malinka et al., 2023; Susnjak, 2022). A recent study (Susnjak, 2022) reveals that ChatGPT is capable of generating logically consistent answers across disciplines, balancing both depth and breadth. Another quantitative analysis (Malinka et al., 2023) shows that students using ChatGPT (by keeping or refining the results from LLMs as their own answers) perform better than average students in some courses from the field of computer security. Recently, several perspective papers (Tan et al., 2023; Kamalov and Gurrib, 2023) also explore various application scenarios of LLMs in classroom teaching, such as teacher-student collaboration, personalized learning, and assessment automation. However, the application of LLMs in education may lead to a series of practical issues, e.g., plagiarism, potential bias in AI-generated content, overreliance on LLMs, and inequitable access for non-English speaking individuals (Kasneci et al., 2023).
To provide researchers with a broad overview of the domain, numerous exploratory and survey papers have been proposed. For example, Qadir (2023); Rahman and Watanobe (2023) and Rahman and Watanobe (2023) conclude the applications of ChatGPT to engineering education by analyzing the responses of ChatGPT to the related pedagogical questions. Jeon and Lee (2023) and Mogavi et al. (2023) collect the opinions from different ChatGPT user groups, e.g., educator, learner, researcher, through in-person interviews, online post replies and user logs, and conclude the practical applications of LLMs in education scenarios. Baidoo-Anu and Owusu Ansah (2023) and Zhang and Tur (2023) focus on the literature review over the published papers and summarize the progress of the area with structured tables. Although the above works have covered a wide range of existing applications of LLMs in education scenarios and provided their long-term visions for future studies, we argue that none of the literature has systematically summarized LLMs for education from a technological perspective. To bridge this gap, this survey aims to provide a comprehensive technological review of LLMs for education, which provides a novel technology-centric taxonomy and summary of existing publicly available datasets and benchmarks. Furthermore, we summarize current challenges as well as further research opportunities, in order to foster innovations and understanding in the dynamic and ever-evolving landscape of LLMs for education. In summary, our contributions lie in the following three main parts:
@@
3, Current challenges and future research directions We discuss current risks and challenges, as well as highlight future research opportunities and directions, urging researchers to dive deeper into this exciting area.
@@ 2.2. Study Assisting
Providing students with timely learning support has been widely recognized as a crucial factor in improving student engagement and learning efficiency during their independent studies (Dewhurst et al., 2000). Due to the limitation of prior algorithms in generating fixed-form responses, many of the existing study-assisting approaches face poor generalization challenges while being implemented in real-world scenarios (König et al., 2023). Fortunately, the appearance of LLMs brings revolutionary changes to this field. Using finetuned LLMs (Ouyang et al., 2022) to generate human-like responses, recent studies in LLM-based educational support have demonstrated promising results. These studies provide real-time assistance to students by helping them solve challenging questions, correcting errors, and offering explanations or hints for areas of confusion.
@@ 2.2.1. Question Solving (QS)
Contributing to the large-scale parameter size of LLMs and the enormous sized and diverse web corpus used during the pre-training phase, LLMs have been proven to be a powerful question zero-shot solver to questions spread from a wide spread of subjects, including math (Yuan et al., 2023; Wu et al., 2023d), law (Bommarito II and Katz, 2022; Cui et al., 2023), medicine (Thirunavukarasu et al., 2023; Liévin et al., 2023), finance (Wu et al., 2023c; Yang et al., 2023), programming (Kazemitabaar et al., 2023; Savelka et al., 2023), language understands(Zhang et al., 2024; Achiam et al., 2023). In addition, to further improve LLM’s problem-solving performance while facing complicated questions, a variety of studies have been actively proposed. For example, Wei et al. (2022) propose the Chain-of-Thought (CoT) prompting method, which guides LLMs to solve a challenging problem by decomposing it into simpler sequential steps. Other works (Sun, 2023; Wang et al., 2023) exploit the strong in-context learning ability of LLMs and propose advanced few-shot demonstration-selection algorithms to improve LLM’s problem-solving performance to general questions. Chen et al. (2022) and Gao et al. (2023a) leverage external programming tools to avoid calculation errors introduced during the textual problem solving process of raw LLMs. Wu et al. (2023a) regard chat-optimized LLMs as powerful agents and design a multi-agent conversation to solve those complicated questions through a collaborative process. Cobbe et al. (2021) and Zhou et al. (2024b) propose the external verifier module to rectify intermediate errors during generation, which improves LLM problem solving performance on challenging math questions. Overall, with the proposition of all these novel designs, the use of LLMs for question solving has achieved impressive progress. Furthermore, students can find high-quality answers to their blocking questions in a timely manner.
@@ 3. Dataset and Benchmark
LLMs revolutionized the field of natural language processing (NLP) by enabling a wide range of text-rich downstream tasks, which leverage the extensive knowledge and linguistic understanding embedded within LLMs to perform specific functions requiring comprehension, generation, or text transformation. Therefore, many datasets and benchmarks are constructed for text-rich educational downstream tasks. The majority of datasets and benchmarks lie in the tasks of question-solving (QS), error correction (EC), question generation (QG), and automatic grading (AG), which cover use cases that benefit different users, subjects, levels, and languages. Some of these datasets mainly benefit the student, while others help the teacher.
@@
Datasets and benchmarks for educational applications vary widely in scope and purpose, targeting different aspects of the educational process, such as student performance data (Ray et al., 2003), text and resource databases (Brooke et al., 2015), online learning data (Ruipérez-Valiente et al., 2022), language learning database (Tiedemann, 2020), education game data (Liu et al., 2020), demographic and socioeconomic data (Cooper et al., 2020), learning management system (LMS) data (Conijn et al., 2016), special education and needs data (Morningstar et al., 2017). Specifically, the question-solving ones (Cobbe et al., 2021; Hendrycks et al., 2020; Huang et al., 2016; Wang et al., 2017; Zhao et al., 2020; Amini et al., 2019; Miao et al., 2021; Lu et al., 2021b; Kim et al., 2018; Lu et al., 2021a; Chen et al., 2023b), take a significant amount as a prevalent task for both Education and NLP fields. In particular, many datasets (Cobbe et al., 2021; Hendrycks et al., 2020; Huang et al., 2016; Wang et al., 2017; Zhao et al., 2020; Amini et al., 2019; Miao et al., 2021; Lu et al., 2021b) are constructed for math question-solving, aiming to provide an abstract expression from a narrative description. Some datasets also take the images (Miao et al., 2021; Kim et al., 2018; Lu et al., 2021a; Kembhavi et al., 2016) and tables (Lu et al., 2021b) into consideration. On the other hand, another bunch of datasets and benchmarks (Kim et al., 2018; Kembhavi et al., 2016; Chen et al., 2023b), are constructed for science textbook question-solving, which requires a comprehensive understanding of the textbook and provide the answer corresponding to the key information in the question. There are also large amounts of datasets and benchmarks constructed for error correction. They are used for foreign language training (Rothe et al., 2021; Ng et al., 2014; Bryant et al., 2019; Tseng et al., 2015; Zhao et al., 2022; Xu et al., 2022b; Du et al., 2023; Náplava et al., 2022; Rozovskaya and Roth, 2019; Grundkiewicz and Junczys-Dowmunt, 2014; Davidson et al., 2020; Syvokon and Nahorna, 2021; Cotet et al., 2020) and computer science programming language training (Just et al., 2014; Le Goues et al., 2015; Lin et al., 2017; Tufano et al., 2019; Li et al., 2022; Guo et al., 2024). The foreign language training datasets and benchmarks contains grammatical errors and spelling errors that are needed to be identified and corrected. The programming training datasets and benchmarks include several code bugs that require sufficient coding understanding for proper correction. On the other hand, there are several datasets and benchmarks for teacher-assisting tasks. (Welbl et al., 2017; Lai et al., 2017; Xu et al., 2022a; Chen et al., 2018; Gong et al., 2022; Hadifar et al., 2023; Liang et al., 2018; Bitew et al., 2022) are constructed for question construction task that aims to evaluate the ability of the LLM for generating the educational questions from a given context. (Yang et al., 2023; Tigina et al., 2023; Blanchard et al., 2013; Stab and Gurevych, 2014) are built for automatic grading the student assignments. We summarize the commonly used publicly available datasets and benchmarks for evaluating LLMs on education applications in Table 1 of the Appendix111While there exist various educational applications, including those that assist with confusion helper, material creation, knowledge tracking, and content personalizing, we did not discuss them in this section due to the absence of publicly accessible datasets..
4. Risks and Potential Challenges
4.1. Fairness and Inclusiveness
4.2. Reliability and Safety
4.3. Transparency and Accountability
LLM, by design, runs as a black-box mechanism, so it comes with transparency and accountability concerns. Milano et al. (Milano et al., 2023a) and BaHammam et al. (BaHammam et al., 2023) raise several challenges about LLMs to higher education, including plagiarism, inaccurate reporting, cheating in exams, as well as several other operational, financial, pedagogical issues. As a further thought of students using generative AI in homework or exams, Macneil et al. (MacNeil et al., 2024) discuss the respective impact to the traditional assessment methods, and think we educators should come up with new assessment framework to take into account the usage of Chat-GPT-like tools. Zhou et al. (Zhou et al., 2024a) brought up academic integrity ethical concerns that specifically confused teachers and students and calls for rethinking of policy-making. As specific measures, Gao et al. (Gao et al., 2024a) introduce a novel concept called mixcase, representing a hybrid text form involving both machine-generated and human-generated and developed detectors that can distinguish human and machine texts. To tackle LLM ethical concerns in intellectual property violation, Huang et al. (Huang and Chang, 2023) propose incorporating citation while training LLMs, which could help enhance content transparency and verifiability. Finlayson, et al. (Finlayson et al., 2024) developed a systematic framework to efficiently discover the LLM’s hidden size, obtaining full-vocabulary outputs, detecting and disambiguating different model updates, which could help users hold providers accountable by tracking model changes, thus enhancing the accountability.
4.4. Privacy and Security
4.5. Overly Dependency on LLMs
Given the impressive performance of LLM generative ability, there is great concern that students blindly rely on LLMs for most of their work, leaving their ability to think independently disappearing. Milano et al. (Milano et al., 2023a) discuss the problem of overreliance caused by Chat-GPT-like applications that students may use to compose essays and academic publications without improving their writing skills, which is essential to cultivate critical thinking. The concern might have even more impact on foreign-language students or students who are educationally disadvantaged while placing less emphasis on learning how to craft well-written texts. Krupp et al. (Krupp et al., 2023) discuss the challenges of overreliance on the implications of LLMs in education and propose some moderated approaches to mitigate such effects. Similarly, Zuber et al. (Zuber and Gogoll, 2023) discuss the risk that overreliance can bring to democracy and suggest cultivating thinking skills in children, fostering coherent thought formulation, and distinguishing between machine-generated output and genuine. They argue that LLMs should be used to augment but not substitute human thinking capacities. Adewumi et al. (Adewumi et al., 2023) also present scenarios that students tend to rely on LLMs for essay writing rather than writing their own, and demonstrates that using a probing-chain-of-thought tool could substantially stimulate critical thinking in the accompany with LLMs.
5. Future Directions
Here, we discuss the future opportunities for LLMs in education and summarize the promising directions in Figure 3. For each direction, we discuss the potential application of advanced LLM-based techniques and conclude their impacts on the future of education.
5.1. Pedagogical Interest Aligned LLMs
Although advanced LLMs like GPT-4 have demonstrated promising performance while being applied to experimental applications in education, applying LLMs directly in real-world instruction is still challenging, as delivering high-quality education is a complicated task that involves multi-discipline knowledge and administration constraints (Milano et al., 2023b). To solve these issues, future researchers in this direction could leverage advanced generation techniques like Retrieval-Augmented Generation (RAG) (Gao et al., 2023b) to inform LLMs with the necessary prior information and guide LLMs to generate the pedagogical interest-aligned results. Apart from that, collecting large-size pedagogical instruction datasets from real-world instruction scenarios and fine-tuning existing LLMs to align with human instructors’ behavior will also be an interesting direction to explore for future research. With learning from the preference of human instructors, LLMs could encoder the pedagogical restriction and knowledge patterns within LLMs’ parameter space, and generate pedagogical interest-aligned results without too much intervention from the external information.
5.2. LLM-Multi-Agent Education Systems
The generous usage of LLMs in language comprehensive, reasoning, planning, and programming inspires works like AutoGen (Wu et al., 2023a) in developing a collaboration framework that involves multiple LLMs to solve complicated tasks through conversation-format procedures. Similarly, problems in education commonly involve multi-step processing logic, which will be a good fit for using the multi-agent-based LLMs system. The recent work by Yang et al. (2024) has demonstrated the great potential of the multi-agent framework for grading tasks. In this work, a human-like grading procedure is achieved by leveraging multiple LLM-based grader agents and critic agents, and the discrepancy errors made by individual judges are corrected through the group discussion procedure. For future research in this direction, more types of LLM-based agents can be included and their functions may spread from specific command executors to high-level plan makers. More importantly, human instructors, who are viewed as the special agents in the system, can also get involved with LLMs’ interaction directly and provide any necessary interventions to the system flexibly.
5.3. Multimodal and Multilingual Supports
The high similarity between different human languages naturally enables LLMs to effectively support multilingual tasks. In addition to that, the recent findings of the alignment between multimodal and language tokens (Wu et al., 2023b) further extend LLMs beyond textual analysis, venturing into multimodal learning analytics. By accepting diverse inputs, LLMs can take advantage of the mutual information between different data resources and provide advanced support to challenging tasks in education. For future works in multimodal direction, more attention could be placed on developing LLMs capable of interpreting and integrating these varied data sources, offering more nuanced insights into student engagement, comprehension, and learning styles. Such advances could pave the way for highly personalized and adaptive learning experiences that are tailored to the unique needs of each student. On the other hand, multilingual LLMs provide convenient access to quality global education resources for every individual with his/her proficient language. With developing robust models that not only translate, but also understand cultural nuances, colloquial expressions, and regional educational standards, researches in this direction would help learners around the world to benefit from LLMs in their native languages and significantly improve equity and inclusion in global education.
5.4. Edge Computing and Efficiency
5.5. Efficient Training of Specialized Models
5.6. Ethical and Privacy Considerations
6. Conclusion
Appendix A Appendix
Table 1 summarizes the commonly used publicly available datasets and benchmarks for evaluating LLMs on education applications.